find the coordinates of the centre of a circle and its radius if the equation is 1) x2+y2-10x-12y+6=0

2) -2x2-2y2-18y+9=0


1
Expert's answer
2021-06-15T18:00:59-0400

(1)

"x^2+y^2-10x-12y+6=0"

Now by completing the squares

"x^2-10x+(\frac{10}{2})^2+y^2-12y+(\frac{12}{2})^2+6=(\frac{10}{2})^2+(\frac{12}{2})^2\\\implies(x^2-10x+25)+(y^2-12y+36)=2+36-6\\\implies(x-5)^2+(y-6)^2=55..........................(1)"

standard form of the circle is

"(x-h)^2+(y-k)^2=r^2...................................(2)"

where (h,k)is the center and r is radius of the circle by comparing ed. (1) and (2)

we get,

centre(h,k)=(5,6)

"Radius r=\sqrt{55}"


(2)

"-2x^2-2y^2-18y+9=0\\\implies-2(x^2+y^2+9y-\frac{9}{2})=0\\\implies x^2+y^2 +9y-\frac{9}{2}=0"

by completing squares,

"x^2+y^2+9y+(\frac{9}{2})^2=\frac{9}{2}+(\frac{9}{2})^2\\\implies x^2 +y^2+9y+\frac{81}{4}=\frac{9}{2}+\frac{81}{4}\\\implies x^2+(y+\frac{9}{2})^2=\frac{18+81}{4}\\\implies(x-0)^2+(y+(-\frac{9}{2}))^2=\frac{99}{4}"


By comparing from standard form of circle we get,

centre (h,k) ="(0,\frac{-9}{2})"


Radius "r=\sqrt{\frac{99}{4}}=\frac{3}{4}\sqrt11"